# Partial domain adaptation enables cross domain cell type annotation between scRNA-seq and snRNA-seq

### v1

@@ Abstract
Accurate cell type annotation across datasets is a key challenge in single-cell analysis. snRNA-seq enables profiling of frozen or difficult-to-dissociate tissues, complementing scRNA-seq by capturing fragile or rare cell types. However, cross-annotation between these two datasets remains largely unexplored, as existing methods treat them independently. We introduce ScNucAdapt, the first method designed for cross-annotation between scRNA-seq and snRNA-seq datasets. To address distributional and cell composition differences, ScNucAdapt employs partial domain adaptation. Experiments across diverse samples show that ScNucAdapt achieves robust and accurate cell type annotation, outperforming existing approaches. Therefore, ScNucAdapt provides a practical framework for the cross-domain cell type annotation between scRNA-seq and snRNA-seq data.
@@ 1 Introduction
This design enables ScNucAdapt to focus on cell types that are shared across datasets while minimizing the negative impact of non-overlapping or dataset-specific cell types. Moreover, ScJoint[scjoint], ScNCL[yan2023scncl], ScCobra[zhao2025sccobra], and ScCorrect[liu2025sccorrect], which utilize transfer learning and domain adaptation methods for the label transfer from unpaired ScRNA-seq to ScATAC-seq datasets, also gave us inspiration on developing ScNucAdapt.
To the best of our knowledge, ScNucAdapt is the first method to focus on cross-annotation between ScRNA-seq and SnRNA-seq datasets. In summary, the contributions of our proposed method can be listed as follows:
@@
    ScNucAdapt is the first cross-domain annotation framework that enables robust label transfer between scRNA-seq and snRNA-seq.
@@
The following passages are organized as follows. In Section 2, the methods of ScNucAdapt are presented, and the descriptions of the datasets used and the evaluation metrics are included. In Section 3, the experimental results on the classification accuracy between a set of ScRNA-seq and SnRNA-seq are presented, including an ablation experiment showing the effectiveness of each component in ScNucAdapt, and a sensitivity analysis. In Section 4, the discussion is presented. Finally, in Section 5, the conclusion is presented.
2 Methods and Datasets
@@ 3.3 ScNucAdapt Enables Cross-Annotation of Mouse Cortical Cell Types Across scRNA-seq and snRNA-seq
The experimental results shown in Table 5 revealed that the cross-annotation from scRNA-seq to snRNA-seq, ScNucAdapt, achieves an accuracy score of 99.78, which performed better than SingleCellNet, ScMap, and scAdapt. On the cross-annotation from snRNA-seq to ScRNA-seq, ScNucAdapt, ScAdapt and SingleCellNet achieved an accuracy of 100.00, while ScMap achieved 99.13.
@@ 3.4 Ablation Experiments
To test these hypotheses, ablation experiments were performed on five cross-domain cell-type annotation tasks, including datasets on bladder cell types, kidney cell types, and tumor cell types. The results shown in Figure 4 reveal that excluding either component substantially decreases generalization performance, indicating that both CS-based distributional alignment and dynamic cluster selection are critical for robust cross-domain annotation. In particular, removing the dynamic cluster selection module, while retaining the CS divergence, still led to a noticeable decline in accuracy, underscoring the complementary role of adaptive clustering in enhancing ScNucAdapt’s ability to generalize between scRNA-seq and snRNA-seq domains. These findings highlight the importance of both proposed components in mitigating modality-specific distributional differences and achieving stable cross-domain cell-type classification.
@@ 4 Discussion
While ScNucAdapt is the first method proposed for the cross-domain annotation between scRNA-seq and snRNA-seq and shows robustness, several problems and questions could be addressed in future work.
@@ 5 Conclusion
In this work, we introduced ScNucAdapt, a novel framework designed for cross-domain cell-type annotation between scRNA-seq and snRNA-seq datasets. To the best of our knowledge, ScNucAdapt is the first method to explicitly address this challenge by integrating principles from partial domain adaptation with dynamic clustering and distribution alignment through Cauchy–Schwarz divergence. This design enables ScNucAdapt to selectively transfer relevant information from annotated data to unlabeled data, mitigating the impact of non-overlapping cell types and modality-specific biases.
@@ Acknowledgments
Figure 1: Concept of Partial domain adaptation
Figure 2: Overall Framework of ScNucAdapt
Figure 3: Concept of Dynamic selection of Clusters in Target data
[…]

### v2

@@ Abstract
Accurate cell type annotation across datasets is a key challenge in single-cell analysis. snRNA-seq enables profiling of frozen or difficult-to-dissociate tissues, complementing scRNA-seq by capturing fragile or rare cell types. However, cross-annotation between these two datasets remains largely unexplored, as existing methods treat them independently. We introduce ScNucAdapt, a method designed for cross-annotation between scRNA-seq and snRNA-seq datasets. To address distributional and cell composition differences, ScNucAdapt employs partial domain adaptation. Experiments across both unpaired and paired scRNA-seq and SnRNA-seq shows that ScNucAdapt achieves robust and accurate cell type annotation, outperforming existing approaches. Therefore, ScNucAdapt provides a practical framework for the cross-domain cell type annotation between scRNA-seq and snRNA-seq data.
@@ 1 Introduction
Figure 1: Concept of Partial domain adaptation
This design enables ScNucAdapt to focus on cell types that are shared across datasets while minimizing the negative impact of non-overlapping or dataset-specific cell types. Moreover, ScJoint[scjoint], ScNCL[yan2023scncl], ScCobra[zhao2025sccobra], and ScCorrect[liu2025sccorrect], which utilize transfer learning and domain adaptation methods for label transfer from unpaired ScRNA-seq to ScATAC-seq datasets, also provided inspiration for developing ScNucAdapt.
To the best of our knowledge, Our Study is the first to focus on cross-annotation between paired or unpaired ScRNA-seq and SnRNA-seq datasets. In summary, the contributions of our proposed method can be listed as follows:
@@
    ScNucAdapt is a cross-domain annotation framework that enables robust label transfer between paired or unpaired scRNA-seq and snRNA-seq datasets.
@@
Figure 2: Overall Framework of ScNucAdapt
The following passages are organized as follows. In Section 2, the methods of ScNucAdapt are presented, and the descriptions of the datasets used and the evaluation metrics are included. In Section 3, the experimental results on classification accuracy between a set of ScRNA-seq and SnRNA-seq are presented, including an ablation experiment that demonstrates the effectiveness of each component in ScNucAdapt, as well as a sensitivity analysis. In Section 4, the discussion is presented. Finally, in Section 5, the conclusion is presented.
2 Materials and Method
@@ 2.2 Dynamic Selection of Clusters in Target data
Figure 3: Concept of Dynamic selection of Clusters in Target data
@@ 3.3 ScNucAdapt Enables Cross-Annotation of Mouse Cortical Cell Types Across scRNA-seq and snRNA-seq
The experimental results shown in Table 5 revealed that the cross-annotation from scRNA-seq to snRNA-seq, ScNucAdapt, achieves an accuracy score of 99.78, which performed better than SingleCellNet, ScMap, and scAdapt. On the cross-annotation from snRNA-seq to ScRNA-seq, ScNucAdapt, ScAdapt, and SingleCellNet achieved an accuracy of 100.00, while ScMap achieved 99.13.
@@ 3.4 Ablation Experiments
To test these hypotheses, ablation experiments were performed on five cross-domain cell-type annotation tasks, including datasets on bladder cell types, kidney cell types, and tumor cell types. The results shown in Figure 4 reveal that excluding either component substantially decreases generalization performance, indicating that both CS-based distributional alignment and dynamic cluster selection are critical for robust cross-domain annotation.
[…]
In particular, removing the dynamic cluster selection module, while retaining the CS divergence, still led to a noticeable decline in accuracy, underscoring the complementary role of adaptive clustering in enhancing ScNucAdapt’s ability to generalize between scRNA-seq and snRNA-seq domains. These findings highlight the importance of both proposed components in mitigating modality-specific distributional differences and achieving stable cross-domain cell-type classification.
@@ 4 Discussion
While ScNucAdapt is proposed for the cross-domain annotation between scRNA-seq and snRNA-seq and shows robustness, several problems and questions could be addressed in future work.
@@ 5 Conclusion
In this work, we introduced ScNucAdapt, a novel framework designed for cross-domain cell-type annotation between scRNA-seq and snRNA-seq datasets. To the best of our knowledge, ScNucAdapt is a method to explicitly address this challenge by integrating principles from partial domain adaptation with dynamic clustering and distribution alignment through Cauchy–Schwarz divergence. This design enables ScNucAdapt to selectively transfer relevant information from annotated data to unlabeled data, mitigating the impact of non-overlapping cell types and modality-specific biases.